115V incandescent lamps A and B are connected in series across a 230V


source. If lamp A is rated 75watts and lamp B is rated 50watts, determine


the current drawn by the series connection.

1
Expert's answer
2022-02-25T02:52:02-0500

A is rated 75watts and lamp B is rated 50watts

P1= V1I1= 75=115*I1

I1= 1.533 A

R1= 115/1.533 = 75.016 ohm

P2= V2I2= 50=115*I2

I2= 0.434 A

R2= 115/0.434=264.976 ohm

Total Resistance =339.976

Current in circuit = 230/ 339.976= 0.67666 A